a person travelling in a straight line moves witha constant velocity v1 for certain distance x and with a constant velocity v2 for next equal distance. the average velocity v is given by the relation?

Grade:12th pass

1 Answers

Vibekjyoti Sahoo
2 years ago
During first phase, time taken=x/v1
during second phase time taken=x/v2
Total distance=2x
Total time taken= x/v1+x/v2
=x[1/v1+1/v2]
Average velocity=2x/x[1/v1+1/v2]v=2/1/v1+1/v22/v=1/v1+1/v2
Hope it helped.